could you please fill in the blank for me, 1,5,13,25,_,61,85
Answer by Iluvidina(15)

The answer is 41
The pattern is adding each number by the next multiple of 4
1+4=5
5+8=13
13+12=25
25+16=41
41+20=61
and
61+24=85